DEAR SIR , 1 . The minute hand of a clock is 7cm long . find the displacement and distance covered by its npointed end when it moves from 6. 00 A . M to 6.30 A .M .

DEAR  SIR  ,  
 
    1 .  The minute hand of a clock is 7cm long . find the displacement and distance covered by its npointed end when it moves from 6. 00 A . M  to  6.30 A .M . 

Grade:9

2 Answers

Aziz Alam IIT Roorkee
9 years ago
At 6 am minute hand is pointing at upward position and at 6:30, it is downward.
The two points are two opposite ends of diameter of circle. So, the displacement is 2*R = 14 cm.
Distance covered is half of the circumference i.e. = pi*R = 3.14*7 =21.98 cm


Thanks and Regards
askIITians Faculty
Sourabh Singh IIT Patna
9 years ago

In one hour minute hand completes one round of clock i.e. 3600. So 6 to 6:30 is half an hour which means 360/2 i.e 1800.
So the distance will be Half the circumference of the circle having radius 7 is 22*7/7= 22cm
and the displacement will be the diameter i.e. 2*7=14 cm